OpenStudy (anonymous):

Simplify the expression. 3 log 3 (6x – 4)

OpenStudy (anonymous):

6x-4, I think. 3log3=1 (3 to what power equals 3) distribute the 1, and you lose the parenthesis can't do much more than that, unless u want to factor out a 2, which doesn't change anything

OpenStudy (anonymous):

it probably means write as \[\log_3((6x-4)^3)\] but there is nothing "simpler" about either one

OpenStudy (anonymous):

3log3(6x-4) 3log18x-3log4 when you subtract logs means you are dividing 3log18x=3log12 X=3log12/log18 =12/18 =2/3

OpenStudy (anonymous):

what the heck????

OpenStudy (anonymous):

forget mine... I did it wrong.

OpenStudy (anonymous):

first answer is wrong. for example \[3\log_3(10)=\log_3(10^3)\] which sure as hell is not 10

OpenStudy (anonymous):

second answer makes no sense at all. you cannot distribute the base. nor can you break apart the log as \[log(x-y)=log(x)-log(y)\]

OpenStudy (anonymous):

3log3(10)=log3(10^3) this true